a nail of mass m1 kg is being hammered by a hammer of mass m2 kg with a velocity of v m/s such that nail drives by s meter into the wall, find the average resistance offered by the wall on the nail???pl answer

Grade:10

1 Answers

Naveen Kumar
9 years ago
Let the hammer transfer all its momentum to the nail during the process.
so applying conservation of momentum,
m2*v=m1*v’
v’=m2*v/m1
Hence initial kinetic energy of the nail= (1/2)*m1*v’2= (1/2)*m22*v2*1/m1=H(let)
Final kinetic energy=0
Now let F is the average force applied by the wall on the nail.So energy E=-F*s
acoording to energy conservation principle this energy must be equalto change in kinetic energy of the nail=H=E
hence avg resistance,F=E/s=H/s